Can you solve these problems by substitution?
1.)
x + 2y = -1
4x - 5y = 22
2.)
-3x + 7y = 14
2x - y = -13

Answer by Maths68(1474)


1.)
x + 2y = -1............(1)
4x - 5y = 22...........(2)
Multiply (1) by -4 and add to (2)
-4x-8y=4
4x-5y=22
-13y=26
-13y/-13=26/-13
y=-2
Put the value of y in (1)
x+2y=-1
x+2(-2)=-1
x-4=-1
x=-1+4
x=3
x,y = 3, -2
Check
Put the values of x and y in (1)
x + 2y = -1............(1)
3+2(-2)=-1
3-4=-1
-1=-1

2.)
-3x + 7y = 14 ..........(1)
2x - y = -13...........(2)
Multiply (2) by 7 and add to (2)
-3x + 7y = 14
14x - 7y = -91
--------------
11x=-77
11x/11=-77/11
x=-7
Put the value of x in (2)
2x - y = -13...........(2)
2(-7)-y=-13
-14-y=-13
-y=-13+14
-y=1
y=-1
x,y=-7,-1
Check
Put the values of x and y in (2)
2(-7) - (-1) = -13
-14 + 1 = -13
-13=-13